What is the speed of light in water?
1
Expert's answer
2020-11-27T14:07:09-0500

We can find the speed of light in water from the Snell's Law:


"\dfrac{v_2}{v_1}=\dfrac{n_1}{n_2},"

here, "v_1=3\cdot10^8\ \dfrac{m}{s}" is the speed of light in the air, "v_2" is the speed of light in water, "n_1=1.0" is the index of refraction of the air and "n_2=1.33" is the index of refraction of water.

Then, from this formula we can find the speed of light in water:


"v_2=\dfrac{v_1n_1}{n_2}=\dfrac{3\cdot10^8\ \dfrac{m}{s}\cdot 1.0}{1.33}=2.25\cdot10^8 \ \dfrac{m}{s}."

Answer:

"v_2=2.25\cdot10^8 \ \dfrac{m}{s}."